# GFM-RAG: Graph Foundation Model for Retrieval Augmented Generation

### v1

@@ 3.2.2 Training Process
${\mathcal{L}}_{\text{BCE}}=-\frac{1}{|{\mathcal{A}}_{q}|}\sum_{e\in{\mathcal{A}}_{q}}\log P_{q}(e)-\frac{1}{|{\mathcal{E}}^{\texttt{-}}|}\sum_{|{\mathcal{E}}^{\texttt{-}}|}\log(1-P_{q}(e)),$
where ${\mathcal{A}}_{q}$ denotes the set of target relevant entities to the query $q$, and -${\mathcal{E}}^{\texttt{-}}\subseteq{\mathcal{E}}\setminus{\mathcal{A}}_{q}$ denotes the set of negative entities sampled from the KG. However, due to the sparsity of the target entities, the BCE loss may suffer from the gradient vanishing problem (Lin et al., 2024). To address this issue, we further introduce the ranking loss (Bai et al., 2023) to maximize the margin between the positive and negative entities:
@@
Unsupervised KG Completion Pre-training. To enhance the graph reasoning capability of the GFM retriever, we first pre-train it on a large-scale knowledge graph (KG) completion task. We sample a set of triples from the KG index and mask either the head or tail entity to unsupervisedly create synthetic queries in the form $q=(e,r,?)~{}\text{or}~{}(?,r,e^{\prime})$, with the masked entity serving as the target entity ${\mathcal{A}}_{q}=\{e\}~{}\text{or}~{}\{e^{\prime}\}$. The GFM retriever is then trained to predict the masked entity using both the query and the KG, as outlined in equation 13.
Supervised Document Retrieval Fine-tuning. After unsupervised pre-training, we fine-tune the GFM retriever on a supervised document retrieval task. In this task, queries $q$ are natural language questions, and target entities ${\mathcal{A}}_{q}$ are extracted from labeled supporting documents ${\mathcal{D}}_{q}$. The GFM retriever is trained to retrieve relevant entities from the KG index using the same training objective as in equation 13.
@@ 3.3 Documents Ranking and Answer Generation
These retrieved entities are then used by the document ranker to obtain the final documents. To diminish the influence of popular entities, we weigh the entities by the inverse of their frequency as entities mentioned in the document inverted index $M\in\{0,1\}^{|{\mathcal{E}}|\times|{\mathcal{D}}|}$ and calculate the final document relevance scores by summing the weights of entity mentioned in documents:
@@ 4 Experiment
In experiments, we aim to address the following research questions: (1) How does GFM-RAG perform in multi-hop retrieval and QA tasks? (Sections 4.2 and 4.3); (2) What are the efficiency and effectiveness of GFM-RAG in multi-hop retrieval? (Section 4.4); (3) How well does GFM-RAG generalize to unseen datasets as a foundation model? (Section 4.6); (4) How does the performance of GFM-RAG scale with training as a foundation model? (Section 4.7); (5) How to interpret GFM-RAG in multi-hop reasoning? (Section 4.8).
@@ 4.1 Experimental Setup
Datasets. We first evaluate the effectiveness of GFM-RAG on three widely-used multi-hop QA datasets, including HotpotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2022), and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020). For a fair comparison (Trivedi et al., 2023; Gutiérrez et al., 2024), we use 1,000 samples from each validation set for testing. We merge the candidate passages into the document corpus and extract 20,000 samples from each training set for GFM training. We also evaluate the performance of GFM-RAG on seven RAG datasets from three domains, including biomedical (Jin et al., 2019), custom support (Sadat et al., 2023; Nandy et al., 2021; Malaviya et al., 2023; Castelli et al., 2020), and general knowledge (Nguyen et al., 2016; Kamalloo et al., 2023), to demonstrate the generalizability of GFM-RAG as the foundation model. The detailed statistics of the test datasets are shown in the Appendix A.
Baselines. We compare against several widely used retrieval methods under three categories: (1) single-step naive methods: BM25 (Robertson & Walker, 1994), Contriever (Izacard et al., 2022), GTR (Ni et al., 2022), ColBERTv2 (Santhanam et al., 2022), RAPTOR (Sarthi et al., 2024), Proposition (Chen et al., 2024); (2) state-of-the-art graph-enhanced methods: LightRAG (Guo et al., 2024), HippoRAG (Gutiérrez et al., 2024); (3) multi-step methods: IRCoT (Trivedi et al., 2023), which can be integrated with arbitrary retrieval methods to conduct multi-step retrieval and reasoning. The detailed introduction of the baselines are shown in the Appendix B.
Metrics. For retrieval performance, we use recall@2 (R@2) and recall@5 (R@5) as evaluation metrics. For the final QA performance, we use the EM score and F1 score following previous works (Gutiérrez et al., 2024).
Implementation Details. The GFM retriever is implemented with 6 query-dependent message passing layers with the hidden dimension set to 512. The pre-trained all-mpnet-v2 (SBERT, 2021) is adopted as the sentence embedding model and is frozen during training. The total parameters of the GFM retriever are 8M, which is trained on 8 NVIDIA A100s (80G) with batch size 4, learning rate 5e-4, and loss weight $\alpha=0.3$. The training data contains 60 KGs with over 14M triples constructed from 700k documents extracted from the training set. The statistics are shown in Table 1, and the detailed data construction process, model settings, and training process are shown in Appendix C.
Table 1: Statistics of the query-doc pairs and KGs used for training.
Dataset #Q-doc Pair #Document #KG #Entity #Relation #Triple
HotpotQA 20,000 204,822 20 1,930,362 967,218 6,393,342
MuSiQue 20,000 410,380 20 1,544,966 900,338 4,848,715
2Wiki 20,000 122,108 20 916,907 372,554 2,883,006
Total 60,000 737,310 60 4,392,235 2,240,110 14,125,063
@@
Multi-step IRCoT + BM25 65.6 79.0 34.2 44.7 61.2 75.6
@@ 4.2 Retrieval Performance
We first evaluate the retrieval performance of GFM-RAG against the baselines on three multi-hop QA datasets. As shown in Table 2, GFM-RAG achieves the best performance on all datasets, outperforming the SOTA IRCoT + HippoRAG by 16.8%, 8.3%, 19.8% in R@2 on HotpotQA, MuSiQue, and 2Wiki, respectively. The results demonstrate the effectiveness of GFM-RAG in multi-hop retrieval. From the result, we can observe that the naive single-step retrievers (e.g., BM25, RAPTOR) are outperformed by graph-enhanced HippoRAG, which highlights the significance of graph structure in multi-hop retrieval. Although LightRAG uses the graph structure, it struggles with multi-hop QA tasks because its retriever lacks multi-hop reasoning capability. With the help of LLMs, the multi-step retrieval pipeline IRCoT improves the performance of all single-step methods through iterative reasoning and retrieval. However, GFM-RAG still outperforms the multi-step IRCoT + HippoRAG by a large margin even with a single-step retrieval. This indicates that the GFM-RAG can effectively conduct the multi-hop reasoning in a single step (detailed in Section 4.8 and Section D.6), which is more efficient and effective than the multi-step retrieval pipeline (detailed in Section 4.4).
@@
Multi-step IRCoT (ColBERTv2) 45.5 58.4 19.1 30.5 35.4 45.1
@@
4.3 Question Answering Performance
We then evaluate the QA performance of GFM-RAG, as it is directly influenced by retrieval quality. We adopt the GPT-4o-mini (OpenAI, 2024a) as LLM and use the top-$5$ retrieved documents for generating answers. From the results shown in Table 3, the single-step GFM-RAG has already achieved state-of-the-art performance against all other baselines. Meanwhile, we also integrate GFM-RAG with IRCoT to conduct multi-step retrieval and reasoning, which further improves the performance by 8.5%, 21.2%, 3.9% in EM on three datasets, respectively. The results demonstrate the effectiveness and great compatibility of GFM-RAG with an arbitrary multi-step framework in multi-hop reasoning tasks.
@@ 4.3 Question Answering Performance
4.4 Efficiency Analysis
GFM-RAG achieves great efficiency in performing multi-step reasoning in a single step. As shown in Table 4, while the naive single-step methods get the best efficiency whose performance is not satisfying. Admittedly, the multi-step framework IRCoT could improve the performance, but it suffers from high computational costs due to the iterative retrieval and reasoning with LLMs. In contrast, GFM-RAG conducts multi-hop reasoning within a single-step GNN reasoning, which is more effective than single-step methods and more efficient than multi-step ones.
@@ 4.5 Ablation Study
We conduct ablation studies to investigate the effectiveness of different components in GFM-RAG, including different sentence embedding models (Section D.1), pre-training strategies (Section D.2), and loss weighting strategies (Section D.3). The results show that GFM-RAG is not sensitive to different sentence embedding models, and the pre-training strategy, as well as the loss weighting strategy, are both crucial for the performance of GFM-RAG.
Figure 3: Model performance and generalizability comparison.
@@ 4.6 Model Generalizability
To demonstrate the generalizability of GFM-RAG as a foundation model, we test the performance (R@5) of GFM-RAG on seven domain-specific RAG datasets without any fine-tuning. Specifically, we first build the KG-index from the documents in each dataset. Then, given the query, we use the pre-trained GFM retriever to retrieve the top-$K$ documents with the help of the corresponding KG-index. As shown in Figure 3, GFM-RAG achieves the best performance on all datasets, outperforming the SOTA HippoRAG by 18.9% on average. The results demonstrate the generalizability of GFM-RAG as the foundation model which can be directly applied to various unseen datasets without any fine-tuning. Additionally, results in Section D.4 demonstrate GFM-RAG’s strong transferability for further performance improvement when fine-tuned on domain-specific datasets.
@@
Question What football club was owned by the singer of ”Grow Some Funk of Your Own”?
@@ 4.7 Model Neural Scaling Law
We further investigate the neural scaling law of GFM-RAG, which quantifies how model performance grows with the scale of training data and model parameter size. It has been validated in the recent foundation models (Kaplan et al., 2020; Dehghani et al., 2023). As shown in Figure 4, the performance of GFM-RAG (MRR: $z$) scales well with the training data ($x$) and the model size ($y$), which can be fitted by the power-law scaling law $z\propto 0.24x^{0.05}+0.11y^{0.03}$. The results demonstrate the scalability of GFM-RAG as the foundation model and potential for further improvement. The detailed analysis of the neural scaling law is shown in Section D.5.
@@ 4.8 Path Interpretations
GFM-RAG exhibits the multi-hop reasoning ability powered by the multi-layer GFM. We provide path interpretations of GFM-RAG for multi-hop reasoning in Table 5. Inspired by NBFNet (Zhu et al., 2021), the paths’ importance to the final prediction can be quantified by the partial derivative of the prediction score with respect to the triples at each layer (hop), defined as:
@@
The top-$k$ path interpretations can be obtained by the top-$k$ longest paths with beam search. We illustrate the path interpretations in Table 5. In the first example, GFM-RAG successfully deduces that the singer of the song has a football club named after him and that he owned it. In the second example, GFM-RAG identifies two paths related to the murder case and the judge presiding over the trial. These interpretations show that GFM-RAG exhibits the ability of multi-hop reasoning within single-step retrieval. We also illustrate the distribution the multi-hop prediction in Section D.6.
@@ 5 Conclusion
Impact Statement
@@
[…]
Appendix A Datasets
A.1 Multi-hop QA Datasets
Three multi-hop QA datasets are used in our experiments: HotpotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2022), and 2WikiMultiHopQA (2Wiki) (Ho et al., 2020). We provide a brief overview of these datasets below.
  • HotpotQA (Yang et al., 2018) is a multi-hop QA dataset that requires reasoning over multiple documents to answer questions. The dataset consists of 97k question-answer pairs, where each question is associated with up to 2 supporting and several distracting documents. The questions are designed to be answerable using multiple pieces of information from the supporting documents.
  • MuSiQue (Trivedi et al., 2022) is a challenging multi-hop QA dataset with 25k 2-4 hop questions. It requires coherent multi-step reasoning to answer questions that span multiple documents.
  • 2WikiMultiHopQA (2Wiki) (Ho et al., 2020) is a multi-hop QA dataset that requires reasoning over multiple Wikipedia articles to answer questions. The dataset consists of 192k questions, which are designed to be answerable using information from 2 or 4 articles.
In experiments, we follow existing methods (Trivedi et al., 2023; Gutiérrez et al., 2024) to use the same 1,000 samples from each validation set and merge the candidate passages as the document corpus for KG-index construction, whose statistics are presented in Table 6.
A.2 Domain-specific RAG Datasets
To test the generalizability of GFM-RAG, we evaluate it on seven domain-specific RAG datasets (Friel et al., 2024) including, (1) biomedical: PubMedQA (Jin et al., 2019); (2) customer support: DelucionQA (Sadat et al., 2023), TechQA (Castelli et al., 2020), ExpertQA (Malaviya et al., 2023), EManual (Nandy et al., 2021); (3) general knowledge: MS Marco (Nguyen et al., 2016), HAGRID (Kamalloo et al., 2023). We provide a brief overview of these datasets below.
  • PubMedQA (Jin et al., 2019) is a collection of PubMed research abstracts with corresponding questions paired with 4 abstract chunks.
  • DelucionQA (Sadat et al., 2023) is a domain-specific RAG dataset leveraging Jeep’s 2023 Gladiator model manual as the source of knowledge, where each question is associated with 4 context documents and only 1 relevant passage.
  • TechQA (Castelli et al., 2020) is a collection of real-world user questions posted on IBMDeveloper and DeveloperWorks forums, along with 10 technical support documents relating to each question.
  • ExpertQA (Malaviya et al., 2023) is a collection of curated questions from domain experts in various fields of science, arts, and law. The dataset also contains expert-curated passages relevant to each question.
  • EManual (Nandy et al., 2021) is a question-answering dataset comprising consumer electronic device manuals and realistic questions about them composed by human annotators, where each question is related with up to 3 context documents.
  • MS Marco (Nguyen et al., 2016) is an open-domain question-answering dataset sourced from Bing search engine user query logs. Each question is associated with 10 context passages retrieved via Bing web search.
  • HAGRID (Kamalloo et al., 2023) is a multi-lingual information retrieval dataset with questions and passages from MIRACL (Zhang et al., 2022).
In experiments, we use test sets constructed by RAGBench (Friel et al., 2024) and merge all the candidate passages as document corpus for KG-index construction. The statistics of the test dataset are detailed in Table 6.

### v2

@@ 3.2.2 Training Process
${\mathcal{L}}_{\text{BCE}}=-\frac{1}{|{\mathcal{A}}_{q}|}\sum_{e\in{\mathcal{A}}_{q}}\log P_{q}(e)-\frac{1}{|{\mathcal{E}}^{\texttt{-}}|}\sum_{|{\mathcal{E}}^{\texttt{-}}|}\log(1-P_{q}(e)),$ (11)
where ${\mathcal{A}}_{q}$ denotes the set of target relevant entities to the query $q$, and -${\mathcal{E}}^{\texttt{-}}\subseteq{\mathcal{E}}\setminus{\mathcal{A}}_{q}$ denotes the set of negative entities sampled from the KG. However, due to the sparsity of the target entities, the BCE loss may suffer from the gradient vanishing problem [36]. To address this issue, we further introduce the ranking loss [2] to maximize the margin between the positive and negative entities:
@@
Self-supervised KG Completion Pre-training. To enhance the graph reasoning capability of the GFM retriever, we first pre-train it on a large-scale knowledge graph (KG) completion task. We sample a set of triples from the KG index and mask either the head or tail entity to create synthetic queries in the form $q=(e,r,?)~\text{or}~(?,r,e^{\prime})$, with the masked entity serving as the target entity ${\mathcal{A}}_{q}=\{e\}~\text{or}~\{e^{\prime}\}$. The GFM retriever is then trained to predict the masked entity using both the query and the KG, as outlined in equation 13.
Supervised Document Retrieval Fine-tuning. After self-supervised pre-training, we supervised fine-tune the GFM retriever on a labeled document retrieval task. In this task, queries $q$ are natural language questions, and target entities ${\mathcal{A}}_{q}$ are extracted from labeled supporting documents ${\mathcal{D}}_{q}$. The GFM retriever is trained to retrieve relevant entities from the KG index using the same training objective as in equation 13.
@@ 3.3 Documents Ranking and Answer Generation
These retrieved entities are then used by the document ranker to obtain the final documents. To diminish the influence of popular entities, we weight the entities by the inverse of their frequency as entities mentioned in the document inverted index $M\in\{0,1\}^{|{\mathcal{E}}|\times|{\mathcal{D}}|}$ and calculate the final document relevance scores by summing the weights of entity mentioned in documents:
@@ 4 Experiment
In experiments, we aim to address the following research questions: (1) How does GFM-RAG perform in multi-hop retrieval and QA tasks? (Sections˜4.2 and 4.3); (2) What are the efficiency and effectiveness of GFM-RAG in multi-hop retrieval? (Section˜4.4); (3) How well does GFM-RAG generalize to unseen datasets as a foundation model? (Section˜4.6); (4) How does the performance of GFM-RAG scale with training as a foundation model? (Section˜4.7); (5) How to interpret GFM-RAG in multi-hop reasoning? (Section˜4.8).
@@ 4.1 Experimental Setup
Datasets. We first evaluate the effectiveness of GFM-RAG on three widely-used multi-hop QA datasets, including HotpotQA [72], MuSiQue [63], and 2WikiMultiHopQA (2Wiki) [20]. We also evaluate the performance of GFM-RAG on seven RAG datasets from three domains, including biomedical [25], custom support [54, 44, 39, 4], and general knowledge [45, 27], to demonstrate the generalizability of GFM-RAG as the foundation model. The detailed statistics of the test datasets are shown in the Appendix˜B.
Baselines. We compare against several widely used retrieval methods under three categories: (1) single-step naive methods: BM25 [53], Contriever [22], GTR [46], ColBERTv2 [55], RAPTOR [56], Proposition [6]; (2) graph-enhanced methods: GraphRAG (MS) [9], LightRAG [15], HippoRAG [16], SubgraphRAG [32], G-retriever [18]; (3) multi-step methods: Adaptive-RAG [23], FLARE [24], and IRCoT [64] framework that can be integrated with arbitrary retrieval methods to conduct multi-step retrieval and reasoning. The detailed introduction of the baselines are shown in the Appendix˜C.
Metrics. For retrieval performance, we use recall@2 (R@2) and recall@5 (R@5) as evaluation metrics. For final QA performance, we use the EM score and F1 score following previous works [16].
Implementation Details. The GFM retriever is implemented with 6 query-dependent message passing layers with the hidden dimension set to 512. The pre-trained all-mpnet-v2 [57] is adopted as the sentence embedding model and is frozen during training. The total parameters of the GFM retriever are 8M, which is trained on 8 NVIDIA A100s (80G) with batch size 4, learning rate 5e-4, and loss weight $\alpha=0.3$. The training data contains 60 KGs with over 14M triples constructed from 700k documents extracted from the training set. The statistics of training data are shown in Table˜5, and the implementations are detailed in Appendix˜D.
@@
SubgraphRAG 61.5 73.0 42.1 49.3 70.7 85.5
G-retriever 53.3 65.5 38.8 45.1 60.8 67.8
Multi-step Adaptive-RAG 61.0 76.4 35.1 44.7 44.7 61.4
FLARE 73.1 81.3 44.3 55.1 67.1 73.1
IRCoT + BM25 65.6 79.0 34.2 44.7 61.2 75.6
@@ 4.2 Retrieval Performance
We first evaluate the retrieval performance of GFM-RAG against the baselines on three multi-hop QA datasets. As shown in Table˜1, GFM-RAG achieves the best performance on all datasets, outperforming the SOTA IRCoT + HippoRAG by 16.8%, 8.3%, 19.8% in R@2 on HotpotQA, MuSiQue, and 2Wiki, respectively. The results demonstrate the effectiveness of GFM-RAG in multi-hop retrieval. From the result, we can observe that the naive single-step retrievers (e.g., BM25, RAPTOR) are outperformed by graph-enhanced HippoRAG, which highlights the significance of graph structure in multi-hop retrieval. Although GraphRAG (MS) and LightRAG use the graph structure, it struggles with multi-hop QA tasks as its retriever is designed for summarization and lacks multi-hop reasoning capability. With the help of LLMs, the multi-step retrieval pipeline IRCoT improves the performance of all single-step methods through iterative reasoning and retrieval. However, GFM-RAG still outperforms the multi-step methods by a large margin even with a single-step retrieval. This indicates that the GFM-RAG can effectively conduct the multi-hop reasoning in a single step (detailed in Section˜4.8 and Section˜E.8), which is more efficient and effective than the multi-step retrieval pipeline (detailed in Section˜4.4).
4.3 Question Answering Performance
@@
Multi-step Adaptive-RAG 45.5 59.6 13.8 25.6 48.9 62.8
FLARE 48.7 60.6 16.2 28.4 46.7 65.4
IRCoT (ColBERTv2) 45.5 58.4 19.1 30.5 35.4 45.1
@@
We then evaluate the QA performance of GFM-RAG, as it is directly influenced by retrieval quality. We adopt the GPT-4o-mini [47] as LLM and use the top-$5$ retrieved documents for generating answers. From the results shown in Table˜2, the single-step GFM-RAG has already achieved state-of-the-art performance against all other baselines. Meanwhile, we also integrate GFM-RAG with IRCoT to conduct multi-step retrieval and reasoning, which further improves the performance by 8.5%, 21.2%, 3.9% in EM on three datasets, respectively. The results demonstrate the effectiveness and great compatibility of GFM-RAG with an arbitrary multi-step framework in multi-hop reasoning tasks.
4.4 Efficiency Analysis
@@
GFM-RAG achieves great efficiency in performing multi-step reasoning in a single step. As shown in Table˜3, while the naive single-step methods get the best efficiency whose performance is not satisfying. Admittedly, the multi-step framework IRCoT could improve the performance, but it suffers from high computational costs due to the iterative retrieval and reasoning with LLMs. In contrast, GFM-RAG conducts multi-hop reasoning within a single-step GNN reasoning, which is more effective than single-step methods and more efficient than multi-step ones.
@@ 4.5 Ablation Study
We conduct ablation studies to investigate the effectiveness of different components in GFM-RAG, including: different sentence embedding models (Section˜E.1), pre-training strategies (Section˜E.2), loss weighting strategies (Section˜E.3), ranking methods (Section˜E.4), training datasets (Section˜E.5), and the construction of KG-index (Section˜E.9). The results show that GFM-RAG is not sensitive to different sentence embedding models, and the pre-training strategy, as well as the loss weighting strategy, are both crucial for the performance of GFM-RAG.
@@ 4.6 Model Generalizability
To demonstrate the generalizability of GFM-RAG as a foundation model, we test the performance (R@5) of GFM-RAG on seven RAG datasets without any domain-specific fine-tuning. Specifically, we first build the KG-index from the documents in each dataset. Then, given the query, we use the pre-trained GFM retriever to retrieve the top-$K$ documents with the help of the corresponding KG-index. As shown in Figure˜4, GFM-RAG achieves the best performance on all datasets, outperforming the SOTA HippoRAG by 18.9% on average. The results demonstrate the generalizability of GFM-RAG as the foundation model which can be directly applied to various unseen datasets without any domain-specific fine-tuning. Additionally, results in Section˜E.6 demonstrate GFM-RAG’s strong transferability for further performance improvement when fine-tuned on domain-specific datasets.
Figure 3: Model generalizability comparison.
@@
Question What football club was owned by the singer of "Grow Some Funk of Your Own"?
@@ 4.7 Model Neural Scaling Law
We further investigate the neural scaling law of GFM-RAG, which quantifies how model performance grows with the scale of training data and model parameter size. It has been validated in the recent foundation models [29, 7]. As shown in Figure˜4, the performance of GFM-RAG (MRR: $z$) scales well with the training data ($x$) and the model size ($y$), which can be fitted by the power-law scaling law $z\propto 0.24x^{0.05}+0.11y^{0.03}$. The results demonstrate the scalability of GFM-RAG as the foundation model and potential for further improvement. The detailed analysis of the neural scaling law is shown in Section˜E.7.
@@ 4.8 Path Interpretations
GFM-RAG exhibits the multi-hop reasoning ability powered by the multi-layer GFM. We provide path interpretations of GFM-RAG for multi-hop reasoning in Table˜4. Inspired by NBFNet [78], the paths’ importance to the final prediction can be quantified by the partial derivative of the prediction score with respect to the triples at each layer (hop), defined as:
@@
The top-$k$ path interpretations can be obtained by the top-$k$ longest paths with beam search. We illustrate the path interpretations in Table˜4. In the first example, GFM-RAG successfully deduces that the singer of the song has a football club named after him and that he owned it. In the second example, GFM-RAG identifies two paths related to the murder case and the judge presiding over the trial. These interpretations show that GFM-RAG exhibits the ability of multi-hop reasoning within single-step retrieval. We also illustrate the distribution the multi-hop prediction in Section˜E.8.
@@ 5 Conclusion
Acknowledgments
G Haffari is partly supported by the ARC Future Fellowship FT190100039 and DARPA Assured Neuro Symbolic Learning and Reasoning (ANSR) program under award number FA8750-23-2-1016. C Gong is supported by NSF of China (Nos: 62336003, 12371510). D Phung is supported by the Australian Research Council (ARC) Discovery Project DP250100262 and DP230101176. S Pan was partly funded by Australian Research Council (ARC) under grants FT210100097 and DP240101547 and the CSIRO – National Science Foundation (US) AI Research Collaboration Program.
@@ Acknowledgments
Appendix A Query-dependent GNNs for Multi-hop Reasoning and Retrieval
We provide a detailed explanation about why query-dependent GNNs can be used for multi-hop reasoning and retrieval. Recent studies [21, 52] have theoretically proven that query-dependent GNNs are effective for capturing the multi-hop logical associations on KGs to answer queries, such as:
$\exists y:\texttt{politician\_of}(\text{Barack Obama},y)\leftarrow\texttt{work\_in}(\text{Barack Obama},z_{1})\wedge\texttt{city\_of}(z_{1},y),$ (20)
where the right part denotes the logical associations can be executed to answer the query on the left, i.e., “politician_of(Barack Obama,y)”.
This query is semantic equivalent to the nature language question: “Barack Obama is the politician of which country?”. By treating the input question as a “soft query” (query in nature language), we apply the query-dependent GNN (GFM) to bridge the gap between nature language and logical query. The GFM tries to understand the semantic of the questions and learn to conduct complex logical reasoning (e.g., multi-hop reasoning) on KGs for retrieval [73]. The learned logical associations for reasoning are shown in Section˜4.8.
Table 5: Statistics of the query-doc pairs and KGs used for training.
Dataset #Q-doc Pair #Document #KG #Entity #Relation #Triple
HotpotQA 20,000 204,822 20 1,930,362 967,218 6,393,342
MuSiQue 20,000 410,380 20 1,544,966 900,338 4,848,715
2Wiki 20,000 122,108 20 916,907 372,554 2,883,006
Total 60,000 737,310 60 4,392,235 2,240,110 14,125,063